OpenStudy (patrickcharran97):

HELP I SUCK USING SUBSTITUTION! solve using substitution: -4-6y=-6 -8x-6y=-18

OpenStudy (anonymous):

is it -4x in start?

OpenStudy (anonymous):

pick an equation first or second then solve for either x or y

OpenStudy (patrickcharran97):

yes its -4x sorry

OpenStudy (anonymous):

mulitiply eq1 by -1 both side 4x+6y=6 add in equation 2 we get -4x=-12 hence x=3 hence 6y= 6-3*4 =-6 y=-1 solution set(3,-1)